OpenStudy (anonymous):

A student is solving the equation 4^x–1 = 64^x+3. Step 1: 4^x – 1 = 64^x + ^3 Step 2: 4^x – 1 = (43)^x + ^3 Step 3: 4^x – 1 = 43^x + ^9 Which is the next step? 4 = 3x + 9 x - 1 = -(3x + 9) x - 1 = x + 3 x - 1 = 3x + 9

OpenStudy (anonymous):

\[4^{x-1} = 64^{x+3}\]

OpenStudy (rsadhvika):

Step 3: \(\LARGE 4^{x – 1} = 4^{3x + 9} \)

OpenStudy (rsadhvika):

since the bases are equal, exponents also will be equal. So the next step would be to equate the exponents

OpenStudy (anonymous):

\[Step 1: 4^{x-1} = 64^{x+3} \] \[Step 2: 4^{x-1}= (43)^{x+3} \] \[Step 3: 4^{x-1} = 43^{x+9}\]

OpenStudy (rsadhvika):

It has to be like this : \[\large Step 1: 4^{x-1} = 64^{x+3}\] \[\large Step 2: 4^{x-1} = (4^3)^{x+3}\] \[\large Step 3: 4^{x-1} = 4^{3x+9}\] \[\large \color{green}{ Step 4:x-1 =3x+9}\]

OpenStudy (rsadhvika):

64 CANNOT become 43 just like that, okay ?

OpenStudy (rsadhvika):

3 needs to be in the exponent : \(\large 64 = 4\times 4\times 4 = 4^3\)

OpenStudy (triciaal):

where did the 43 come from?

OpenStudy (triciaal):

oh it should be 4^3x

OpenStudy (triciaal):

(4^3)^3 to get 4^9 to raise a power to a power multiply

OpenStudy (triciaal):

step 4 as given by rsdhvika above choice x-1 = 3x + 9

OpenStudy (anonymous):

Thank you!